# LAPO: Internalizing Reasoning Efficiency via Length-Adaptive Policy Optimization

### v1

@@ 1 Introduction
Existing approaches to address this challenge fall into three main categories, each with inherent limitations. Direct length reduction methods either rely on reward design (Yang et al., 2025; Huang et al., 2025) that can cause over-shortening and accuracy degradation, or impose hard length constraints (Aggarwal & Welleck, 2025; Hou et al., 2025) that lack adaptability across problem types. Dynamic early-stopping approaches (Qiao et al., 2025; Muennighoff et al., 2025) make real-time termination decisions but often truncate mid-reasoning, disrupting the thinking process. Adaptive thinking methods (Lou et al., 2025; Zhang et al., 2025; Fang et al., 2025) enable models to switch between thinking and non-thinking modes but operate at a coarse granularity.
@@
We introduce Length-Adaptive Policy Optimization (LAPO), a two-stage reinforcement learning framework that progressively builds this adaptive capability. In the first stage, we design length-aware rewards that encourage efficiency while maintaining accuracy. During this process, we collect statistical patterns from successful solutions, specifically focusing on the reasonable length range where most correct answers naturally fall. This reveals the inherent reasoning requirements of each problem without imposing artificial constraints. In the second stage, we leverage these discovered patterns to provide explicit guidance to the model. By incorporating target length information directly into the problem prompt, we enable models to plan their reasoning trajectory before beginning the solution process. Crucially, to ensure models can reason adaptively without requiring predefined lengths at inference time, we embed the length constraint as a self-declarative statement immediately after the ¡think¿ token. This technique reframes the budget not as an external command, but as part of the model’s own internal reasoning plan. By learning to generate a solution that aligns with its self-proposed budget, the model is incentivized to internalize the link between problem complexity and resource allocation, allowing it to reason flexibly and efficiently when deployed.
@@ 3.2 Discovery Stage: Learning Natural Reasoning Patterns
$[L_{\min},L_{\max}]=[\text{Percentile}_{30}(\mathcal{L}_{q}),\text{Percentile}_{70}(\mathcal{L}_{q})]$ (2)
@@
$R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$ (4)
The length component operates on a crucial principle—only correct responses receive length-based rewards:
$R_{\text{length-1}}(r_{i},q)=\begin{cases}1.0&\text{if correct and }L_{\min}\leq|r_{i}|\leq L_{\max},\\ \max(0,1-d_{i}/100)&\text{if correct and }|r_{i}|\notin[L_{\min},L_{\max}],\\ 0&\text{if incorrect;}\end{cases}$ (5)
where $d_{i}=\min(||r_{i}|-L_{\min}|,||r_{i}|-L_{\max}|)$. This design creates gradients guiding models toward efficient lengths while allowing flexibility for complex problems. Throughout the Discovery stage, we continuously update $\mathcal{M}$ after each GRPO training step to reflect evolving model capabilities.
@@ 3.3 Internalization Stage: Length-Aware Efficient Reasoning
$\text{prompt}_{q}^{\prime}=\text{prompt}_{q}+\text{``{<think>} I will answer the question with }n\text{ tokens.''}$.” (6)
@@
To encourage the model to follow its self-declared reasoning budget from Equation 6, the Internalization stage employs a precision-focused reward function. This function is designed to reward the alignment between the model’s output length and its self-declared budget n. The total reward is defined as:
$R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$ (7)
where the adherence component, $R_{\text{Length-2}}$, is only granted for correct solutions:
$R_{\text{Length-2}}(r_{i},n)=\begin{cases}\exp\left(-\frac{(|r_{i}|-n)^{2}}{2\sigma^{2}}\right)&\text{if correct},\\ 0&\text{otherwise};\end{cases}$ (8)
@@
8:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q)$
@@
10:         Compute range: $[L_{\min},L_{\max}]=[\text{P}_{30}(\mathcal{L}_{q}),\text{P}_{70}(\mathcal{L}_{q})]$
@@
12:         Compute rewards: $R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$
@@
21:         Augment prompt: $q^{\prime}=q+\text{``{<think>} I will answer the question with }\mathcal{M}(q)\text{ tokens.''}$.”
22:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q^{\prime})$
23:         Compute rewards: $R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$
@@
A cornerstone of our framework is how it fosters genuine internalization, enabling inference-time flexibility without explicit length targets. The key lies in the design of the augmented prompt (Equation 6). By placing the length guidance ”I will answer the question with $n$ tokens.” immediately after the <think> token, we transform an external constraint into an intrinsic part of the model’s cognitive process.
@@ 3.4 Training Pipeline
(Lines 4-13): The model explores natural reasoning patterns through GRPO training with length-aware rewards. For each problem in the training batch, we generate multiple rollouts and extract statistics from successful responses. The mapping $\mathcal{M}$ is continuously updated to capture the evolving understanding of appropriate reasoning lengths. This stage runs for $E_{1}$ epochs, allowing the model to discover problem-specific length patterns through self-supervised exploration.
@@
(Lines 15-24): The model learns to internalize efficient reasoning by incorporating discovered length patterns into the training process. Each problem prompt is augmented with target length information derived from the Discovery stage. The placement of this guidance within the <think> block encourages the model to treat the budget as part of its own reasoning plan, which fosters genuine length awareness rather than rote instruction following. The dual-strategy update mechanism refines the mapping $\mathcal{M}$ throughout training, allowing newly solved problems to establish benchmarks while encouraging efficiency improvements for previously solved ones.
@@ 4 Experiment Setup
We train our models on a mixed dataset of 10,000 mathematical problems. This dataset comprises 6,000 examples randomly sampled from the DeepScaleR-Preview-Dataset (Luo et al., 2025b), which sources problems from competitive math contests (e.g., AIME, AMC) and benchmarks (e.g., Omni-Math (Gao et al., 2024), STILL (Min et al., 2024)). The remaining 4,000 examples are drawn from the intermediate levels (3-5) of the MATH dataset (Hendrycks et al., 2020) to ensure a balanced difficulty distribution. For the base models, we experiment with DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned reasoner, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), an advanced RL-finetuned version known for high-quality but verbose outputs.
All models are trained using the Group Relative Policy Optimization (GRPO) algorithm. The LAPO framework is applied in two sequential stages. Stage 1 (LAPO-Discovery) is trained for 240 steps (approx. 3 epochs) with the length-aware reward $R_{1}$ (Eq. 4), where hyperparameter $\alpha$ is 0.7. The resulting model then serves as the initialization for Stage 2 (LAPO-Internalization), which is trained for another 240 steps using the plan-adherence reward $R_{2}$ (Eq. 7) with $\beta$ set to 0.8. Due to computational constraints, the maximum context length is limited to 4,096 tokens during training. More training details are provided in Appendix A.1.
@@
We evaluate the trained models on four challenging benchmarks representing a spectrum of mathematical difficulty: MATH-500 (Hendrycks et al., 2020), AIME2024, AMC23, and Olympiad-Bench (He et al., 2024). Following standard practices (Guo et al., 2025), we use a sampling temperature of 0.6, a top-p of 0.95, and a maximum generation length of 32,768 tokens. For each benchmark, we report Pass@1 accuracy and the average number of tokens per response. The number of samples per query is 4 for MATH-500 and Olympiad-Bench, and 32 for the more competitive AIME2024 and AMC23.
@@
We benchmark LAPO against two classes of baselines: the foundational models we build upon and existing methods designed for efficient reasoning. First, we evaluate the Base Models to establish a performance starting point. These include DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned model, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), a powerful RL-finetuned reasoner known for its high-quality but often verbose outputs. Both are evaluated without any length-control modifications. Second, we compare against state-of-the-art Efficient Reasoning Baselines that represent two distinct technical approaches. For methods focusing on direct length budgeting, we evaluate L1-Exact and L1-Max(Aggarwal & Welleck, 2025). Both train a model to adhere to a user-specified length constraint by penalizing deviations from the target budget, with L1-Exact enforcing a precise match and L1-Max treating the budget as a soft upper limit. For the iterative pruning strategy, we compare against ThinkPrune(Hou et al., 2025). This method trains a model to shorten its reasoning by imposing a hard token limit during its iterative RL process, where any response exceeding the limit receives a zero reward. We evaluate its two publicly available variants, ThinkPrune-iter2k and ThinkPrune-4k.
@@ 5 Results and Analysis
We present comprehensive experimental results and analysis to validate LAPO’s effectiveness and understand its underlying mechanisms. We begin with the main results (Section 5.1), benchmarking LAPO against baselines and state-of-the-art methods. We then conduct in-depth ablation studies on key design choices, including the the form of length guidance (Section 5.2) and the statistical metrics for target length selection (Section 5.3). Finally, we provide a mechanistic analysis of how LAPO works, examining its emergent ability for difficulty-aware resource allocation (Section 5.4), its qualitative refinement of reasoning patterns (Section 5.5).
5.1 Main Results
As shown in Table 1, LAPO consistently improves reasoning efficiency across all benchmarks, achieving substantial token reduction while simultaneously enhancing or maintaining accuracy. Our key findings are summarized below.
Our framework demonstrates clear superiority over both the foundational models and state-of-the-art baselines. When applied to DeepScaleR-1.5B-Preview, LAPO-I achieves a remarkable 37.8% token reduction while boosting the average accuracy by 2.3 percentage points. This trend holds for DeepSeek-R1-1.5B, where it cuts token usage by 40.9% and improves accuracy by 1.2 points. These results validate that LAPO enables models to generate more concise yet effective reasoning chains.
Compared to existing methods, LAPO establishes a more effective accuracy-efficiency frontier. On the MATH-500 benchmark, our LAPO-I model (86.3%) substantially outperforms both L1-Max (81.9%) and L1-Exact (80.6%). While L1-Max achieves aggressive token compression, its over-shortening appears to compromise reasoning quality on complex problems. In contrast, LAPO strikes a more effective balance. Against ThinkPrune-4k, which shares our 4k training window, LAPO demonstrates superior performance in both token efficiency and accuracy, confirming its ability to find a better trade-off.
The effectiveness of LAPO stems from its two-stage design, where performance progresses verbeter from the Discovery stage (LAPO-D) to the Internalization stage (LAPO-I). LAPO-D first establishes a strong foundation, achieving a significant 36.0% token reduction on its own by learning natural reasoning length distributions. Building on this, LAPO-I achieves an additional 6.9% efficiency gain by internalizing these patterns through in-context guidance. This progressive refinement, combined with LAPO’s consistent performance improvements across diverse benchmarks, confirms that our framework learns generalizable principles of adaptive reasoning rather than task-specific heuristics.
Table 1: Main results on MATH500, AIME2024, AMC, and OlympiadBench. We report Pass@1 accuracy (%) and the average number of generated tokens (#Tok). In each base model group, bold indicates the best and underline indicates the second-best Pass@1 score.
@@ 5.1 Main Results
Base model: DeepSeek-R1-distill-Qwen-1.5B
@@
5.2 Ablation Study on the Length Guidance for LAPO-I
To investigate the effectiveness of the in-context length guidance, we conduct an ablation study on the prefill content used in LAPO-I. We compare our default method, Exact Length Guidance (which provides a precise target $n$), against two variants: Range-based Guidance (providing a target range) and Implicit Guidance (no length information, relying only on the reward).
As shown in Table 2, explicit and precise guidance is crucial for performance. Our default method, LAPO-I, achieves the highest average accuracy (64.8%). However, switching to a range-based guidance results in a 0.7% accuracy drop (to 64.1%), and removing the guidance entirely causes a further drop to 63.6%. This experimental results directly validate the effectiveness of the problem-to-length mapping learned in LAPO-D; the more precisely the model is guided to follow these discovered patterns, the better it performs.
Notably, the Implicit Guidance variant not only suffers from lower accuracy but also sees its token consumption revert to a higher 4,088 tokens, nearly identical to the LAPO-D baseline. This demonstrates that the reward signal alone is insufficient to internalize efficient reasoning behaviors. The explicit, in-context guidance acts as a critical bridge, enabling the model to access and leverage the reasoning templates discovered in LAPO-D.

### v2

@@ 1 Introduction
Existing approaches to address this challenge fall into three main categories, each with inherent limitations. Direct length reduction methods either rely on reward design (Yang et al., 2025; Huang et al., 2025a) that can cause over-shortening and accuracy degradation, or impose hard length constraints (Aggarwal & Welleck, 2025; Hou et al., 2025) that lack adaptability across problem types. Dynamic early-stopping approaches (Qiao et al., 2025; Muennighoff et al., 2025) make real-time termination decisions but often truncate mid-reasoning, disrupting the thinking process. Adaptive thinking methods (Lou et al., 2025; Zhang et al., 2025a; Fang et al., 2025a) enable models to switch between thinking and non-thinking modes but operate at a coarse granularity.
@@
We introduce Length-Adaptive Policy Optimization (LAPO), a two-stage reinforcement learning framework that progressively builds this adaptive capability. In the first stage, we design length-aware rewards that encourage efficiency while maintaining accuracy. During this process, we collect statistical patterns from successful solutions, specifically focusing on the reasonable length range where most correct answers naturally fall. This reveals the inherent reasoning requirements of each problem without imposing artificial constraints. In the second stage, we leverage these discovered patterns to provide explicit guidance to the model. By incorporating target length information directly into the problem prompt, we enable models to plan their reasoning trajectory before beginning the solution process. Crucially, to ensure models can reason adaptively without requiring predefined lengths at inference time, we embed the length constraint as a self-declarative statement immediately after the <think> token. This technique reframes the budget not as an external command, but as part of the model’s own internal reasoning plan. By learning to generate a solution that aligns with its self-proposed budget, the model is incentivized to internalize the link between problem complexity and resource allocation, allowing it to reason flexibly and efficiently when deployed.
@@ 3.2 Discovery Stage: Learning Natural Reasoning Patterns
$L_{\min},L_{\max}]=[\text{Percentile}_{30}(\mathcal{L}_{q}),\text{Percentile}_{70}(\mathcal{L}_{q})]$ (2)
@@
$R_{D}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{1}}(r_{i},q)$ (4)
The length component operates on a crucial principle—only correct responses receive length-based rewards. Let $\mathcal{C}_{i}=\mathbb{I}(y_{i}=y_{\text{gold}})$ indicate whether the response is correct, and define the distance to the target length range as $d_{i}=\min(||r_{i}|-L_{\min}|,||r_{i}|-L_{\max}|)$. We introduce a linear decay function $f(d)=\max(0,1-d/100)$ to penalize deviations from the efficient length range. The length reward is then defined as:
$R_{\text{1}}(r_{i},q)=\begin{cases}1.0&\text{if }\mathcal{C}_{i}=1\land|r_{i}|\in[L_{\min},L_{\max}]\\ f(d_{i})&\text{if }\mathcal{C}_{i}=1\land|r_{i}|\notin[L_{\min},L_{\max}]\\ 0&\text{if }\mathcal{C}_{i}=0\end{cases}$ (5)
This design creates gradients guiding models toward efficient lengths while allowing flexibility for complex problems. Throughout the Discovery stage, we continuously update $\mathcal{M}$ after each GRPO training step to reflect evolving model capabilities.
@@ 3.3 Internalization Stage: Length-Aware Efficient Reasoning
$_{q}^{\prime} = prompt$q + “<think> I will answer the question with $n$ tokens.”
@@
To encourage the model to follow its self-declared reasoning budget, the Internalization stage employs a precision-focused reward function. This function is designed to reward the alignment between the model’s output length and its self-declared budget n. The total reward is defined as:
$R_{I}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{2}}(r_{i},q^{\prime})$ (6)
where the adherence component, $R_{\text{2}}$, is only granted for correct solutions:
$R_{\text{2}}(r_{i},n)=\begin{cases}\exp\left(-\frac{(|r_{i}|-n)^{2}}{2\sigma^{2}}\right)&\text{if }\mathcal{C}_{i}=1,\\ 0&\text{if }\mathcal{C}_{i}=0;\end{cases}$ (7)
@@
8:   Generate NN rollouts: $\{r_{1},\dots,r_{N}\}\sim\pi_{\theta}(q)$
@@
10:   Compute range: $L_{\min},L_{\max}]=[\text{P}_{30}(\mathcal{L}_{q}),\text{P}_{70}(\mathcal{L}_{q})]$
@@
12:   $\text{Compute rewards: }R_{D}(r_{i},q)={}\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{1}}(r_{i},q)$
@@
21:   Augment prompt: $q^{\prime}\leftarrow{}q+\text{``{<think>} }\text{I will}$  $\text{ answer the question}\text{ with }\mathcal{M}(q)\text{ tokens.''}$
22:   Generate NN rollouts: $\{r_{1},\dots,r_{N}\}\sim\pi_{\theta}(q^{\prime})$
23:   $\text{Compute rewards: }R_{I}(r_{i},q^{\prime})={}\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{2}}(r_{i},q^{\prime})$
@@
A cornerstone of our framework is how it fosters genuine internalization, enabling inference-time flexibility without explicit length targets. The key lies in the design of the augmented prompt. Placing the self-declarative guidance immediately after the <think> token transforms an external constraint into an intrinsic part of the model’s cognitive plan.
@@ 3.4 Training Pipeline
(Lines 4-15): The model explores natural reasoning patterns through GRPO training with length-aware rewards. For each problem in the training batch, we generate multiple rollouts and extract statistics from successful responses. The mapping $\mathcal{M}$ is continuously updated to capture the evolving understanding of appropriate reasoning lengths. This stage runs for $E_{1}$ epochs, allowing the model to discover problem-specific length patterns through self-supervised exploration.
@@
(Lines 17-27): The model learns to internalize efficient reasoning by incorporating discovered length patterns into the training process. Each problem prompt is augmented with target length information derived from the Discovery stage. The placement of this guidance within the <think> block encourages the model to treat the budget as part of its own reasoning plan, which fosters genuine length awareness rather than rote instruction following. The dual-strategy update mechanism refines the mapping $\mathcal{M}$ throughout training, allowing newly solved problems to establish benchmarks while encouraging efficiency improvements for previously solved ones.
@@ 5 Experiment Setup
We train our models on a mixed dataset of 10,000 mathematical problems to ensure a balanced difficulty distribution, comprising 6,000 examples from the DeepScaleR-Preview-Dataset Luo et al. (2025b) and 4,000 from the intermediate levels of the MATH dataset Hendrycks et al. (2020). We apply LAPO to two base models: DeepSeek-R1-1.5B Guo et al. (2025) and DeepScaleR-1.5B-Preview Luo et al. (2025b).
All models are trained using the Group Relative Policy Optimization (GRPO) algorithm. Discovery Stage(LAPO-D) is trained for 3 episodes with reward $R_{D}$ (Eq. 4), where hyperparameter $\alpha$ is 0.7. The resulting model then serves as the initialization for the subsequent Internalization Stage(LAPO-I), which is trained for 3 episodes using reward $R_{I}$ (Eq. 6) with $\beta$ set to 0.7. These hyperparameters were empirically chosen to provide a strong incentive for efficiency while ensuring correctness remains the primary learning signal. Note that we did not conduct extensive hyperparameter tuning, so one can expect further improvements with additional optimization. Besides, due to computational constraints, the maximum context length is limited to 4,096 tokens during training. Crucially, to ensure a fair comparison, most baselines, including ThinkPrune and L1, were also trained or evaluated under this same 4k context limit. Further hyperparameters are detailed in Appendix.
@@
At inference, we expand the generation window to a generous 32,768 tokens for all models to assess their true, unconstrained reasoning capabilities. This setup allows us to isolate the efficiency gains stemming directly from the LAPO framework, rather than from simple context window limitations. We evaluate on four challenging benchmarks: MATH-500 Hendrycks et al. (2020), AIME2024, AMC23, and Olympiad-Bench He et al. (2024). Following standard practices Guo et al. (2025), we report both Pass@1 accuracy and the average number of tokens. For each problem, we sample N responses (4 for MATH-500/OlympiadBench, 32 for AIME/AMC) with a temperature of 0.6 and a top-p of 0.95.
@@
We benchmark LAPO against three classes of baselines: the foundational models, an ablation baseline, and existing methods designed for efficient reasoning. First, we evaluate the Base Models to establish a performance starting point. These include DeepSeek-R1-1.5B Guo et al. (2025) and DeepScaleR-1.5B-Preview Luo et al. (2025b). Second, to isolate the effect of our length-reward, we also include an Ablation Baseline, denoted as Acc-Only, which is trained with GRPO using only the accuracy reward. Finally, we compare against several state-of-the-art Efficient Reasoning Baselines, which represent different philosophies for achieving efficiency. (1)Implicit Regularization: HAPO Huang et al. (2025b), which uses history-aware rewards. (2)Budget-Driven Control: L1 Aggarwal & Welleck (2025) and ThinkPrune Hou et al. (2025), which follow external length targets. (3)Adaptive Activation: AutoThink Tu et al. (2025), AdaptThink Zhang et al. (2025b), and Thinkless Fang et al. (2025b), which learn a binary think/no-think policy.
@@ 6 Results and Analysis
Table 1: Main results on MATH500, AIME2024, AMC23, and OlympiadBench. We report Pass@1 accuracy (%) and the average number of generated tokens (#Tok). For each metric, bold indicates the best and underline indicates the second-best Pass@1 score within each base model group.
@@
Base model: DeepSeek-R1-1.5B
HAPO 82.2 2288 31.3 8649 67.3 4735 50.1 5024 57.7 5174
AutoThink 83.5 2017 29.7 7084 70.2 3499 51.2 4606 58.6 3825
AdaptThink 81.6 1580 23.9 6432 63.2 2860 48.5 4616 54.3 3871
@@
We present comprehensive experimental results and analysis to validate LAPO’s effectiveness and understand its underlying mechanisms. We begin with the main results (Section 6.1), benchmarking LAPO against baselines and state-of-the-art methods. We then conduct in-depth ablation studies on key design choices, including the the form of length guidance (Section 6.2) and the statistical metrics for target length selection (Section 6.3). And a targeted experiment demonstrating the model’s robust internalization of reasoning efficiency (Section 6.4). Finally, we provide a mechanistic analysis of how LAPO works, examining its emergent ability for difficulty-aware resource allocation (Section 6.5), its qualitative refinement of reasoning patterns (Section 6.6).
6.1 Main Results
As shown in Table 1, LAPO achieves a superior balance of reasoning accuracy and computational efficiency, consistently outperforming its base models and establishing a new state-of-the-art frontier among methods that do not rely on external length controls.
Compared to its base models, LAPO delivers substantial gains. On DeepScaleR-1.5B-Preview, it reduces tokens by 38.5% while boosting average accuracy by 2.3 points; a similar trend holds for DeepSeek-R1-1.5B (41.0% token cut and 1.2 point accuracy gain). This validates that LAPO learns to produce more concise yet effective reasoning.
When compared with leading efficiency methods, LAPO consistently demonstrates a superior accuracy-efficiency trade-off. On the more capable DeepScaleR-1.5B base model, LAPO-I achieves the highest average accuracy among all tested methods. This advantage holds across different baseline paradigms. It surpasses budget-driven methods like ThinkPrune-4k and L1-Max under a fair 4k training context. Compared to implicit regularization methods like HAPO, LAPO shows a clear advantage in preserving accuracy. Furthermore, while adaptive activation methods like AutoThink can be highly token-efficient, they do not reach the same level of reasoning quality. This comprehensive comparison highlights that LAPO’s “Discover-Internalize” process, which fosters an autonomous and continuous length adaptation, leads to a more robust and effective reasoning policy than methods relying on external budgets, progressive compression, or binary mode-switching.
LAPO-D first establishes a strong foundation, achieving a significant 36.0% token reduction on its own by learning natural reasoning length distributions. This is highlighted by comparing it to the Acc-Only baseline. While simply finetuning for accuracy yields some token reduction, LAPO-D’s length-aware reward achieves substantially greater efficiency while also improving average accuracy by 0.5 points. This demonstrates that encouraging conciseness via our reward not only prunes redundant thoughts but also helps the model find more robust reasoning patterns. Building on this superior foundation, LAPO-I achieves an additional 6.9% efficiency gain by internalizing these patterns through in-context guidance. This progressive refinement indicates that our framework learns a generalizable principle of adaptive reasoning.
6.2 Ablation Study on In-Context Guidance
To validate that our method’s success stems from internalizing a self-proposed plan, we ablate the two key factors of our in-context guidance: its form (how precise the guidance is) and its position (whether it’s part of the model’s internal thought process). We compare our default approach (w/ Exact) against three variants: w/ Range (less precise guidance), w/ Outside (placing the guidance before <think>), and w/ Implicit (no guidance, relying only on the reward). As shown in Table 2, the results demonstrate that both form and position are critical for effective internalization.
Our default method outperforms the less precise Range variant, indicating that specific targets discovered in Discovery stage provide a stronger learning signal. More critically, the guidance’s position determines whether the model internalizes a plan or merely follows instructions. Moving the guidance outside the <think> block transforms it into an external command and causes accuracy to drop significantly to 63.9%. This illustrates that the model performs best when the budget is framed as part of its own cognitive plan. Finally, removing the guidance entirely results in the worst performance, with accuracy dropping to 63.6% and token count reverting to the LAPO-D baseline. This indicates that our explicit, properly-positioned, self-declarative guidance is the critical mechanism for internalization.